OpenStudy (anonymous):

A ball is dropped from the top of a building. The height, y, of the ball above the ground (in feet) is given as a function of time, t, (in seconds) by y = 1210 -16^2 (a) Find the velocity of the ball at time t. (b) When does the ball hit the ground, and how fast is it going? Give your answer in feet per second and miles per hour (1 ft/sec = 15/22 mph). Round your answers to two decimal places. t = ? v(t) = ? feet per second v(t) = ? feet per hour

OpenStudy (anonymous):

Well, since I'm learning about derivatives, I assume I need to use that to get my answer. I tried: y = 1210 - 16t^2 = 1210 -32t -1210 = -32t t = - 32.81 I know the sign has to be negative because the height is decreasing but... that's all I got.

ganeshie8 (ganeshie8):

is that for part a ?

OpenStudy (anonymous):

Yeah, I was trying to do part a. I'm guessing I need to apply a for b, so I didn't try that yet.

ganeshie8 (ganeshie8):

part a, is just asking u to get velocity as a function of t; thats all.

ganeshie8 (ganeshie8):

it should be below :- y = 1210 - 16t^2 v = y' = -32t

OpenStudy (anonymous):

Alright, got it! Now I need to figure out how to apply this to (b), which I'm getting wrong as well.

ganeshie8 (ganeshie8):

to apply it to part b, first u need to find time, t, at which ball hits the ground

ganeshie8 (ganeshie8):

any idea how to find that ? (Algebra)

ganeshie8 (ganeshie8):